OpenStudy (anonymous):

homogeneous equation; dy/dx=[2(2y-x)]/x+y; initial value of y(1)=2

OpenStudy (anonymous):

\[\large{\frac{dy}{dx}=\frac{\frac{4y}{x}-\frac{2x}{x}}{\frac{x}{x}+\frac{y}{x}}}\] =\[\frac{dy}{dx}=\frac{4 \frac{y}{x} -2 }{1+\frac{y}{x}}\] let u=y/x y=ux y'=u+u'x \[u+u'x=\frac{4u-2}{1+u}\]\[u'x=\frac{4u-2-u(1+u)}{1+u}\]

OpenStudy (anonymous):

\[u'x=\frac{4u-2-u-u^2}{1+u} = \frac{-u^2+3u-2}{1+u}\]

OpenStudy (anonymous):

ı found c=0 what should I do after?

OpenStudy (anonymous):

now its seprable , after u integrate and find u , you substitute u=y/x in the equation and solve for y

OpenStudy (anonymous):

c? how did u get that?

OpenStudy (anonymous):

u'=du/dx

OpenStudy (anonymous):

separable and we have to integrate

OpenStudy (anonymous):

yeah did u integrate it?

OpenStudy (anonymous):

yes and after ( u-2)^3/(u-1)^2=c/x after u=y/x y=2 and x=1 therefore c=0